OpenStudy (anonymous):

(-23x^4y^0)^-3 / (92x^-2y)^-3

OpenStudy (anonymous):

= (-23)^-3 * x^-12 / ((92)^-3 * x^6 * y^-3) = -(23)^3 * x^-18 / ((23*4)^-3 * y^-3) = -(23)^3 * x^-18 * y^3 / (23)^-3 * 4^(-3) = -(23)^6 * x^(-18) * y^3 / 4^(-3) = -(23)^6 * 4^3 * x^(-18) * y^3 = -(92*23)^3 * x^(-18) * y^3

OpenStudy (anonymous):

scratch that, it's wrong

OpenStudy (anonymous):

= -(23)^-3 * x^-12 / ((92)^-3 * x^6 * y^-3) = -x^-18 / (23^3 * 23^-3 * 4^-3 * y^-3) = -x^18 * 4^3/y^-3 = -64x^18*y^3

OpenStudy (anonymous):

ugh, that's wrong too: = -64x^-18 * y^3

OpenStudy (anonymous):

thats what i got here it is hold on i will copy and paste it

OpenStudy (anonymous):

(-23x^4y^0)^-3 / (92x^-2y^-3) = ( -x^4y^0) / (4x^(-2)y^(-3)^-3 = ((-x^4-^(-2) *y^0-^(-3)) / (4))^-3 = (-x^6y^3) /(4)) = (-x^6)^3 * (y^3)^-3 / 4^-3 = -x^-18 *y^-19 / 1/64 = 64 / -x^18y^9 = (-23x^4y^0)^-3 / (92x^-2y^-3) = 64 / -x^18y^9